OpenStudy (anonymous):

State whether the following sequence is arithmetic; and if so, state the common difference. 18, 3, 1/2, 1/12, ...

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

no? as in there is no answer?

OpenStudy (anonymous):

im not sure about the rules of this but i suppose u could say the common difference is (5/6)n where n is the number before so for 18 it would b 18-(5/6)18=18-15=3 3-(5/6)3=3-5/2=1/2 etc